#1 Argentina Partners is concerned about the possible effects of inflation on its operations. Presently, the company sells 60,000 units for $30 per unit. The variable production costs are $15, and fixed costs amount to $700,000. Production engineers have advised management that they expect unit labor costs to rise by 15 percent and unit materials costs to rise by 10 percent in the coming year. Of the $15 variable costs, 50 percent are from labor and 25 percent are from materials. Variable overhead costs are expected to increase by 20 percent. Sales prices cannot increase more than 10 percent. It is also expected that fixed costs will rise by 5 percent as a result of increased taxes and other miscellaneous fixed charges. The company wishes to maintain the same level of profi t in real dollar terms. It is expected that to accomplish this objective, profi ts must increase by 6 percent during the year. Required a. Compute the volume in units and the dollar sales level necessary to maintain the present profit level, assuming that the maximum price increase is implemented. b. Compute the volume of sales and the dollar sales level necessary to provide the 6 percent increase in profits, assuming that the maximum price increase is implemented. c. If the volume of sales were to remain at 60,000 units, what price increase would be required to attain the 6 percent increase in profits? #2 Alameda Tile sells products to many people remodeling their homes and thinks that they could profitably offer courses on tile installation, which might also increase the demand for their products. The basic installation course has the following (tentative) price and cost characteristics: Tuition . . . . . . . . . . . . . . . . . . . . . . . $ 400 per student Variable costs (tiles, supplies, and so on) . . . . . . 240 per student Fixed costs (advertising, salaries, and so on) . . . . . . . . . . . 80,000 per year Required a. What enrollment will enable Alameda Tile to break even? b. How many students will enable Alameda Tile to make an operating profit of $40,000 for the year? c. Assume that the projected enrollment for the year is 800 students for each of the following (considered independently): 1. What will be the operating profit (for 800 students)? 2. What would be the operating profit if the tuition per student (that is, sales price) decreased by 10 percent? Increased by 20 percent? 3. What would be the operating profit if variable costs per student decreased by 10 percent? Increased by 20 percent? 4. Suppose that fixed costs for the year are 10 percent lower than projected, whereas variable costs per student are 10 percent higher than projected. What would be the operating profit for the year?
Answer #1
Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!
Argentina Partners
Present Income Statement
Sell price                 30.00
Variable production costs                 15.00
Contribution margin                 15.00
Units sold         60,000.00
Contribution       900,000.00
Fixed costs       700,000.00
Operating Profit       200,000.00
Breakup of Variable production costs
Variable production costs                 15.00
50 percent are labor                   7.50 50% of $ 15
25 percent are materials                   3.75 25% of $ 15
25 percent are Variable overhead                   3.75 25% of $ 15
Revised situation Present Increase Revised
Sell price                 30.00 10%                                                       33.00
Labor                   7.50 15%                                                         8.63
Materials                   3.75 10%                                                         4.13
Variable overhead                   3.75 20%                                                         4.50
Contribution margin                 15.00                                                       15.75
Contribution margin ratio 47.73% It is Contribution margin/ Sell Price
Fixed costs       700,000.00 5%                                             735,000.00
Operating Profit       200,000.00 6%                                             212,000.00
Answer to requirement a Answer to requirement b Answer to requirement c
Operating Profit required       200,000.00 A Operating Profit required       212,000.00 H Operating Profit required       212,000.00 H
Fixed costs       735,000.00 B Fixed costs       735,000.00 B Fixed costs       735,000.00 B
Target Contribution       935,000.00 C=A+B Target Contribution       947,000.00 C=H+B Target Contribution       947,000.00 C=H+B
Contribution margin                 15.75 D Contribution margin                 15.75 D Contribution margin ratio 47.73% I
Volume in units         59,365.08 E=C/D Volume in units         60,126.98 E=C/D Dollar sales level 1,984,190.48 J=C/I
Sell price                 33.00 F Sell price                 33.00 F Volume in units         60,000.00 K
Dollar sales level 1,959,047.62 G=E*F Dollar sales level 1,984,190.48 G=E*F Sell price                 33.07 L=J/K
Alameda Tiles
Answer to requirement a Answer to requirement b
Calculation of breakeven students If target profit is $ 40,000.
Tuition fess per student               400.00 Target Profit required         40,000.00
Variable costs per student               240.00 Fixed costs         80,000.00
Contribution margin per student               160.00 Target Contribution       120,000.00
Fixed costs         80,000.00 Contribution margin per student               160.00
Breakeven students               500.00 Number of students               750.00
Answer to requirement c_1
Calculation of operating profits
Tuition fess per student               400.00
Variable costs per student               240.00
Contribution margin per student               160.00
Number of students               800.00
Contribution amount       128,000.00
Fixed costs         80,000.00
Operating profits         48,000.00
Answer to requirement c_2
If Tuition fess decrease by 10% If Tuition fess Increase by 20%
Tuition fess per student               400.00 Tuition fess per student               400.00
Decrease by 10%                 40.00 Increase by 20%                 80.00
Revised Tuition fess per student               360.00 Revised Tuition fess per student               480.00
Variable costs per student               240.00 Variable costs per student               240.00
Contribution margin per student               120.00 Contribution margin per student               240.00
Number of students               800.00 Number of students               800.00
Contribution amount         96,000.00 Contribution amount       192,000.00
Fixed costs         80,000.00 Fixed costs         80,000.00
Operating profits         16,000.00 Operating profits       112,000.00
Answer to requirement c_3
If Variable costs decrease by 10% If Variable costs Increase by 20%
Variable costs per student               240.00 Variable costs per student               240.00
Decrease by 10%                 24.00 Increase by 20%                 48.00
Revised Variable costs per student               216.00 Revised Variable costs per student               288.00
Calculation of operating profits Calculation of operating profits
Tuition fess per student               400.00 Tuition fess per student               400.00
Variable costs per student               216.00 Variable costs per student               288.00
Contribution margin per student               184.00 Contribution margin per student               112.00
Number of students               800.00 Number of students               800.00
Contribution amount       147,200.00 Contribution amount         89,600.00
Fixed costs         80,000.00 Fixed costs         80,000.00
Operating profits         67,200.00 Operating profits           9,600.00
Answer to requirement c_4
If Variable costs Increase by 10% and Fixed cots decrease by 10%
Variable costs per student               240.00
Increase by 10%                 24.00
Revised Variable costs per student               264.00
Fixed costs         80,000.00
Decrease by 10%            8,000.00
Revised Fixed costs         72,000.00
Calculation of operating profits
Tuition fess per student               400.00
Variable costs per student               264.00
Contribution margin per student               136.00
Number of students               800.00
Contribution amount       108,800.00
Fixed costs         72,000.00
Operating profits         36,800.00